OpenStudy (anonymous):

how do I set up this problem ? if the store is offering a 14%discount and yao pays $670.8 what was the marked price?

OpenStudy (anonymous):

does anyone help anyone on here ? or is it just a chat session I have went on here twice and have not gotten a response

OpenStudy (anonymous):

does that mean "after the discount"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

670.80x0.14

OpenStudy (anonymous):

our instructor says we have to show the 5 steps

OpenStudy (anonymous):

if so, then you know \[\$670.80=.86P\] and so \[P=670.80\div .86\]

OpenStudy (anonymous):

so it is A=p-rP

OpenStudy (anonymous):

reasoning that if you get a 14% discount, you pay 86% of the price

OpenStudy (anonymous):

right he said you subtract the 14 from 100

OpenStudy (anonymous):

by the distributive law \[P-rP=(1-r)P\] so no need for the first step just subtract \(.14\) from \(1\) or 14% from 100%

OpenStudy (anonymous):

so it will look like this it would be 670.8-.86=.86P right?